# MENTOR: Guiding Hierarchical Reinforcement Learning with Human Feedback and Dynamic Distance Constraint

### v1

@@ 4.1 Framework formalization
Figure 2: Experimental Domains for Evaluating MENTOR compared with baselines (see Appendix B for additional information).
4.2 Dynamic Distance Constraint
As indicated in Figure 3, humans can provide direction to a target by providing preferences (indicated by the yellow curve) but do not have real-time knowledge of low-level capabilities. Thus multi-level simultaneous learning may be uncoordinated, resulting in subgoals that are too difficult or too simple. To solve this issue, we introduce the DDC. This method limits the range of subgoals based on their distance from the current state (the variation in green shading depicted in the figure), ensuring they are achievable by the low-level policy. This is achieved through a specific formulation:
$\max_{\pi_{hi}}\mathbb{E}[r_{hi}(s,sg,g)]\quad s.t.\quad H(sg,g,k)\leq 0.$ (3)
The function $H(sg,g,k):\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}$ is defined as $\max(d(sg,g)-k,0)$. In this context, $d(sg,g)$ represents the distance between the subgoal $sg$ and the goal $g$. The hyper-parameter $k$ sets the subgoal space range. However, in scenarios like the Four Rooms domain, the Euclidean distance as $d(\cdot)$ may not accurately assess the difficulty of completing a task. For instance, a goal in the top right might be hard to reach despite a low Euclidean distance, indicating the need for a learned step distance measure. Therefore, in GCRL problems, it’s crucial to evaluate the step distance between unreachable subgoals and the current state. If the low-level policy fails to reach a subgoal, assigning a high distance value to these unreached subgoals relative to the current state is essential. Neglecting to do so, high-level policies might incorrectly view challenging subgoals as easy, leading to unrealistic subgoal proposals. To tackle this issue, we recommend incorporating negative sampling into the distance model objective:
$\min_{d}\frac{1}{2}\mathbb{E}_{\tau\sim\rho_{\pi_{lo}},i\in[0,L],j\in[i,L],sg\sim\pi_{hi}(\cdot|s_{0},g)}$ (4)
$\left[(d(s_{i},s_{j})-|i-j|)^{2}]+\mathbbm{1}(s_{L}\neq sg)(d(s_{i},sg)-L)^{2}\right].$
The objective is formulated by minimizing the expected loss across trajectories $\tau$, sampled using the low-level policy $\rho_{\pi_{lo}}$ from recent episodes, and the goals $g$ drawn from the environment’s distribution $\rho_{g}$. $L$ is the length of the trajectory.
In Equation 3, optimizing is challenging because of the strict constraint. To overcome this difficulty, we utilize the relaxation technique, which allows us to establish an unconstrained optimization objective:
$\max_{\pi_{hi}}\mathbb{E}[r_{hi}(s,sg)-\alpha H(sg,g,k)],$ (5)
where $\alpha$ is a balancing coefficient to adjust the influence of reward from human guidance and distance constraint. During the training process, it is necessary to modify the hyper-parameter $k$ to adjust the difficulty of subgoals dynamically. Hence, a fixed coefficient is inadequate to achieve the desired balance. We propose the implementation of a dynamic balancing coefficient which will be updated in real-time to maintain a balance between the rewards and distance constraints.
To effectively incorporate the adjustment into MENTOR, we need to optimize two parameters simultaneously: high-level policy $\theta_{hi}$ and balancing coefficient $\alpha$, converting our maximization into a dual problem:
$\min_{\alpha\geq 0}\max_{\pi_{hi}}\mathbb{E}[r_{hi}(s,sg,g)-\alpha H(sg,g,k)].$ (6)
The distance constraint function $H(\cdot)$ guarantees that the subgoals within a distance $k$. This function treats subgoals that are within the distance $k$ as having zero effects. However, in the case of $\alpha$, we want it to decrease when this constraint does not work, and increase when the high-level policies make decisions that do not satisfy this constraint. Since $H(\cdot)$ is always greater than 0 and the update gradient for $\alpha$ is singular, we need to eliminate the $max(\cdot)$ function from $H(\cdot)$ to achieve automatic updates for $\alpha$. We update the policy $\pi_{hi}$ firstly and then coefficient $\alpha$ following modification for the optimization:
$\pi_{hi}^{*}=\arg\max_{\pi_{hi}}\mathbb{E}[r_{hi}(s,sg,g)-\alpha H(sg,g,k)],$ (7)
$\alpha^{*}$ $=\arg\min_{\alpha}\mathbb{E}[-\alpha\left(d(sg,g)-k\right)].$ (8)
After implementing DDC at the high-level, we test if the low-level can solve the subgoal, modifying the $k$ value to tune the subgoal difficulty, ensuring multiple levels operate in coordination.
5 EXPERIMENTS
@@
In this study, we perform a range of experiments in various domains, as depicted in Figure 2. These domains have been widely used in previous studies [Plappert et al., 2018, Luo et al., 2023, Ghosh et al., 2019]. Through the experiments, we aim to answer the following research questions (RQs):
@@
RQ1: How does the performance of MENTOR compare to baseline models in various domains?
RQ2: What is the impact of Dynamic Distance Constraint within MENTOR?
RQ3: What is the impact of human feedback within MENTOR?
@@
RQ4: What insights can be gained about the individual contributions of key components in the MENTOR framework?
Figure 4: Graphical representation of the success rates for MENTOR in comparison to other baseline methods across different benchmarks on five random seeds. Within the domain of Four rooms, the curve behaves non-smoothly because the starting point and the goal are fixed. Any curves that are not visible in the graph, indicate a zero success rate throughout the trials. These results are aggregated from an average of five individual runs. Additionally, the shaded areas surrounding each curve represent the standard deviation.
5.1 Setup
@@
Benchmarks: We select FetchPush, FetchPickAndPlace, FetchDraw, FetchObsPush, Pusher, and Four rooms as our simulation benchmarks, widely used in research [Andrychowicz et al., 2017, Ghosh et al., 2019, Bougie and Ichise, 2022]. As illustrated, the first five domains involve long-horizon manipulation tasks, while Four rooms focuses on 2D navigation.
@@
Baselines: We have incorporated various learning methods in our baseline implementation, including techniques from HRL, RLHF, RND, dynamic distance learning, and hindsight relabelling. For further details, please see Appendix C.
  • PEBBLE is a RL framework that incorporates human preferences into policy learning [Lee et al., 2021a].
  • HhP applies non-expert human feedback guiding in the high-level structure and RND bonus in the low-level [Bougie and Ichise, 2022]. We implement this baseline together with hindsight relabelling for fast Learning.
  • HER is a classical RL strategy that uses hindsight relabeling Andrychowicz et al. [2017].
  • DDL (Dynamic Distance Learning) learns distance models. The model is utilized to train goal achievement by setting the reward as the negative distance between states and goals. The implementation follows Hartikainen et al. [2019].
5.2 Evaluating MENTOR’s Performance Against Baselines (RQ1)
@@
In our experiments across six domains, we find that MENTOR excels in learning speed and subgoal attainment as evidenced in Figure 4. This assessment uses five random seeds and is evaluated over 50 tests. As can be seen from DDL and PEBBLE curves. DDL and PEBBLE rarely learn effectively in complex GCRL environments (effective in simple domains shown in Section D.1). DDL’s poor performance may be attributed to the absence of guiding signals and the instability of the reward function. Pairwise comparison guidance makes it difficult for PEBBLE to complete subgoal. HER, a classic GCRL algorithm, serves as a benchmark for evaluating other algorithms. Yet, HhP, despite integrating human guidance and curiosity-driven exploration, underperforms compared to HER in FetchPush. Its benefits are also unclear in Pusher and Four Rooms. This indicates HhP’s inadequate use of human guidance and curiosity in exploration. In contrast, MENTOR, by incorporating Dynamic Distance Constraint and Exploration-Exploitation Decoupling, effectively leverages these elements for exploration, achieving faster and more consistent training than DDL, PEBBLE, and HER. Our analysis highlights the superior performance of MENTOR.
5.3 Analyzing the Influence of Dynamic Distance Constraint (RQ2)
@@
Our study will examine DDC’s effects on learning agents, assessing its impact on agent behaviors and the efficiency of its auto-balancing coefficient for algorithm convergence. We also explore DDC’s synergism with human feedback.
@@
Analyzing the impact of automatic balancing coefficient. DDC has a mechanism to automatically adjust the balancing coefficient. The mechanism plays a crucial role in balancing the impact of the reward model and distance constraint. As shown in Figure 6, fixed values of 0.5, 5, and 20 for $\alpha$ are ineffective in learning compared to the automatic adjustment setting. Small alpha values (0.5 and 5) slow DDC convergence. Large alpha (20) enhances learning speed but makes it unstable and sensitive to $\Delta k$. The auto-adjusted $\alpha$ mechanism converges rapidly and doesn’t demand any fine-tuning for $\Delta k$.
Analyzing the overlay effects of DDC and human feedback. Having investigated DDC’s impact, we now turn to assess the combined influence of human guidance and DDC on subgoal formulation. Figure 7 depicts the training phase subgoal distribution, both with and without DDC. To the right of the figure, DDC’s inactivity results in a fragmented learning interaction between levels. While the high-level, aided by human guidance, swiftly navigates to the goal, it neglects the low-level’s execution capabilities, leading to transient and inefficient guidance, evident in the predominance of red subgoals in most regions except the upper-right corner. In contrast, on the figure’s left, with DDC active, the high-level delineates a subgoal sequence starting from the lower right, proceeding to the lower left, then the upper left, and culminating at the upper right corner. This strategy, different from the scenario without DDC, significantly bolsters the efficiency of human guidance, contributing to MENTOR overall effectiveness. We also analyze the heatmap of reward and distance model, presented in Section D.2.
(a) w/ DDC
(b) w/o DDC
@@
5.4 Analyzing the Impact of Human Feedback (RQ3)
Table 1: Cross-analysis on episodes (100 steps per episode) to the success of variables batch queries and query frequency. For each combination of variables, we ran 5 trials on different seeds and recorded the mean and variance.
@@ 5.4 Analyzing the Impact of Human Feedback (RQ3)
Comparing human collected labels and synthetic labels. In our study, we employ two methods for providing human guidance: synthetic labels generated through scripted techniques and human-generated labels. As explained in Figure 8, models trained with human labels exhibited comparable learning rates, with performance differences potentially attributed to noise in human feedback. Comparing experiments with human feedback to those without any guidance suggests that real human feedback is effective for policy training.
Figure 8: Experiments for evaluating MENTOR with script-generated and human-collected labels on the Four rooms domain given 10 labels every 25 episodes. Script-generated labels and non-feedback experiments are based on five random seeds, and the human-collected label experiment is based on one random seed.
5.5 Ablation Studies (RQ4)

### v2

@@ III-C MENTOR Process
Figure 3: Experimental Domains for Evaluating MENTOR compared with baselines. The Pusher and Four Rooms serve as representative domains within the class of Partially Observable Markov Decision Processes (POMDPs), characterized by uncertainties in the state observations which require the algorithm to make decisions based on incomplete information. In contrast, the rest domains, such as FetchObsPush, provide a fully observable state space. For further details, please refer to Appendix A.
IV EXPERIMENTS
@@
In this section, we perform a range of experiments in various commonly used domains [51, 52, 53], as depicted in Figure 3. Through the experiments, we aim to answer the following research questions (RQs):
RQ1
@@
How does the performance of MENTOR in comparison to baseline models in various domains?
RQ2
What is the impact of DDC in MENTOR?
RQ3
What is the impact of human feedback in MENTOR?
RQ4
@@
What insights can be gained about the individual contributions of key components in MENTOR?
IV-A Setup
@@
Benchmarks: We select FetchPush, FetchPickAndPlace, FetchDraw, FetchObsPush, Pusher, and Four rooms as our simulation benchmarks, widely used in research [4, 18, 53]. As illustrated in Figure 3, the first five domains involve long-horizon manipulation tasks, while Four rooms focuses on 2D navigation. In our experiments utilizing the RLHF, we employed two distinct approaches to acquire human labels: manual labeling by humans and synthetic labeling through scripts. The specifics of these label acquisition processes are delineated in the Appendix B. Across the experimental trials, we primarily utilized synthetic labels, with the exceptions explicitly noted.
Hardware: The computer’s hardware is equipped with an AMD Ryzen 7 5700X processor, 32GB of RAM, and an NVIDIA GeForce RTX 3070 Ti graphics card.
@@
Baselines: We have incorporated various learning methods in our baseline implementation, including techniques from HRL, RLHF, RND, dynamic distance learning, and hindsight relabelling.
  • Reinforcement Learning from Human Feedback (RLHF) firstly establishes a reward model based on human feedback and then using it to optimize policy in reinforcement learning. PEBBLE is a RLHF framework that that combines unsupervised pre-training and preference-based learning to significantly improve the sample and feedback efficiency of human-in-the-loop reinforcement learning [32]. The key concept of PEBBLE is to learn a reward model on human preference. Unlike the learning strategy of the reward model in the MENTOR, it has an unsupervised pretraining step via intrinsic reward. Here, we apply the state entropy $H(s)=-\mathbb{E}\log p(s)$. By converting the policy to a simpler version, the pretraining reward function is set as $\text{log}(||s_{i}-s_{i}^{k}||)$ in the batch. This implies that maximizing the distance between a state and its nearest neighbor increases the overall state entropy.
  • Hindsight Relabeling is a data augmentation method for multi-task reinforcement learning that facilitates data sharing across tasks by relabeling experiences to improve sample efficiency. Our baseline, HER [4] is a classical hindsight relabeling strategy. HER enables efficient learning from limited and non-diverse rewards by re-framing failed attempts as successes towards different goals. In our implementation, we utilize a policy to incorporate HER. When sampling batch transitions from the replay buffer, we employ hindsight technology to modify certain parts of the transitions. It is important to mention that the high-level policy differs from the low-level policy of MENTOR in certain aspects. Unlike the low-level policy, the high-level policy solely receives goals from the environment and does not incorporate an RND bonus.
  • Hierarchical Reinforcement Learning with Human Feedback (HRL-HF) integrates human feedback into the generation of high-level subgoals. We follow the architecture of HhP [18]. HhP introduces an HRL method that integrates human preferences and curiosity-driven exploration. By using human preferences to guide decision-making at high-levels and curiosity to promote environmental exploration at low-levels. HhP is considered to be inefficient in combining HRL and RLHF, and it also introduces bias at the low-levels. When we trained HhP, we found that this algorithm was difficult to converge, and to ensure the effectiveness of training, we introduced the hindsight relabelling technique in this algorithm as well.
  • Distance Learning (DL) learns distance models. The model is utilized to train goal achievement by setting the reward as the negative distance between states and goals. This baseline follows DDL [47]. DDL calculates dynamical distances, defining them as the expected number of steps between states in a system. This method leverages unsupervised interactions for learning distances. In the distance evaluation step, the aim is to estimate the dynamic distance between pairs of states visited by a given policy. A distance function, denoted as $d(s_{i},s_{j})$, is utilized for this purpose. To calculate the distance, multiple trajectories, denoted as $\tau$, are generated by rolling out the policy. Each trajectory has a length of T. The empirical distance between states $s_{i}$ and $s_{j}$ in the trajectory $\tau$ is then calculated, where $0\leq i\leq j\leq T$. The empirical distance is simply given by the difference between j and i. To learn the distance function, supervised regression can be employed by minimizing the objective:
    $\mathcal{L}_{d}=\frac{1}{2}\mathbb{E}_{\tau\sim p_{\tau}}\left[\sum_{i\in[0,T]}\sum_{j\in[i,T]}\left(d_{\psi}(s_{i},s_{j})-(j-i)\right)^{2}\right]$.
Figure 4: Graphical representation of the success rates for MENTOR in comparison to other baseline methods across different benchmarks on five random seeds. The shaded areas surrounding each curve represent the standard deviation. Within the Four Rooms domain, the performance curve exhibits non-smooth behavior due to the fixed positions of the starting point and the goal. Consequently, the success rate can abruptly transition from 0% to 100%, leading to the curve with large variance. Any curves that are not visible in the graph, indicate a zero success rate throughout the trials. These results are aggregated from an average of five individual runs.
IV-B Performance Evaluation (RQ1)
@@
In our experiments across six domains, we find that MENTOR excels in learning speed and subgoal attainment as evidenced in Figure 4. This assessment uses five random seeds and is evaluated over 50 tests. As can be seen from DDL and PEBBLE curves. DDL and PEBBLE rarely learn effectively in complex GCRL environments. DDL’s poor performance may be attributed to the absence of guiding signals and the instability of the reward function. Pairwise comparison guidance makes it difficult for PEBBLE to complete subgoal. HER, a classic GCRL algorithm, serves as a benchmark for evaluating other algorithms. Yet, HhP, despite integrating human guidance and curiosity-driven exploration, underperforms compared to HER in FetchPush. Its benefits are also unclear in Pusher and Four Rooms. This indicates HhP’s inadequate use of human guidance and curiosity in exploration. In contrast, MENTOR, by incorporating Dynamic Distance Constraint and Exploration-Exploitation Decoupling, effectively leverages these elements for utilization of human feedback and exploration, achieving faster and more consistent training than DDL, PEBBLE, and HER. Our analysis highlights the superior performance of MENTOR.
TABLE I: Model parameter count and time consumption for the first 1000 episodes on FetchPush domain. Here, ’M’ denotes ’million’ for the quantity of model parameters, and ’min’ denotes ’minutes’ for the duration of time spent.
Algorithm MENTOR HER
low-level actor&critic 0.560M 0.280M
high-level actor&critic 0.189M 0.189M
reward 0.140M -
rnd 0.278M -
distance 0.068M -
total count 1.235M 0.469M
time consumption 18.63$\pm$ 0.05min 10.76$\pm$ 0.08min
However, MENTOR demonstrates superior performance and is equipped with a greater number of components. To investigate the resource consumption of our algorithm, we conducted additional experiments, meticulously recording the model parameter count and computation time for 1000 iterations of MENOTR and HER in the FetchPush environment. As detailed in Table I, our results reveal that MENTOR has approximately 263% more parameters and requires roughly 176% more running time compared to HER. Although our model has a larger number of components, which significantly increases the parameter count, considering that modern GPUs typically have 8GB or more of memory, these models do not demand excessively high computational resources. Moreover, by examining the convergence of HER in Figure 4 and the computational time consumed by both MENTOR and HER algorithms over 1000 episodes, we found that our algorithm consumes 76% more time. However, the convergence speed of our algorithm is more than 3 times that of HER. Therefore, our algorithm remains highly efficient in terms of resource consumption.
IV-C Impact of DDC (RQ2)
@@
Our study will examine DDC’s effects on learning agents, assessing its impact on agent behaviors and the efficacy of three designs in achieving stable algorithmic convergence. We also explore DDC’s synergism with human feedback.
@@
Analyzing the impact of automatic balancing coefficient. DDC has a mechanism to automatically adjust the balancing coefficient. The mechanism plays a crucial role in balancing the impact of the reward model and distance constraint. As shown in Figure 6, fixed values of 0.5, 5, and 20 for $\alpha$ are ineffective in learning compared to the automatic adjustment setting. Small $\alpha$ values (0.5 and 5) slow DDC convergence. Large $\alpha$ (20) enhances learning speed but makes it unstable and sensitive to $\Delta k$.
To gain more insight into the details of difficulty adjustment under different $\alpha$, we recorded the change curve of $k$ ($\Delta k$ of 0.05) under different $\alpha$ settings in Figure 7. Analyzing the $k$ value change curves for different $\alpha$ values reveals how subgoal difficulty adapts throughout training. These curves track the incremental changes in the $k$ value ($\Delta k$ = 0.05) and highlight the influence of the $\alpha$ parameter on the stability and progression of the learning trajectory. Lower $\alpha$ values lead to a more gradual increase in the $k$ value, promoting a stable but slower learning process. In contrast, higher $\alpha$ values cause more pronounced fluctuations, which may accelerate learning but also introduce greater risk of instability. The Auto-$\alpha$ setting demonstrates both rapid and stable $k$ value growth, indicating an adaptive strategy that balances speed and stability in the learning process. These $k$ value curves also serve as indicators of the low-level policy’s progress: consistent increases suggest ongoing improvement, while plateaus or declines may signal the need to reassess the strategy. These analyses highlight the crucial role of the $\alpha$ parameter in shaping the learning process.
Figure 7: The variation curve of $k$ value throughout the training process under different $\alpha$ value and $\Delta k$ is 0.05.
Figure 8: Near-policy sampling and exploratory mechanism effects on reward model in the Four rooms domain. On the top, the oracle reward function in the Four rooms domain indicates where the agent likely earns higher rewards, guiding it from the lower right to the upper right corner. We use the oracle reward to generate synthetic labels to train the reward model following Equation 5. The heatmaps below the oracle reward reflect how the reward model adjusts in response to the agent’s exploration and policy updates within the environment. The heatmaps of reward model for the top and bottom exhibit changes across 2,000 episodes, whereas the heatmap for the center demonstrates alterations throughout 10,000 episodes.
Analyzing the near-policy sampling for reward model training and the exploratory mechanism at the high-level. During the training process of MENTOR, we observe some phenomena that the high-level policy cannot be well guided by the rewarding model. These observations can be ascribed to two factors: (1) the reward model’s inadequate guidance of the current model, and (2) the increase in the $k$ value leading to the agents in local optima, thereby diminishing further exploration of potential subgoals. In Figure 8, the reward heatmap, shaped by human feedback, approximates the oracle function. We observe that when near-policy samples are absent, the reward model tends to emphasize the entirety of the state space. This, in turn, makes it more challenging and computing resource-consuming for the reward model to effectively learn.
Additionally, we have observed that there are instances where the high-level policy becomes stuck in local optima as we incrementally increase the value of $k$ for a long time. This phenomenon is attributed to two reasons: (1) it lacks pretraining at the high-level, and (2) the high-level lacks the exploratory ability, leading to not trying new subgoals. Pretraining plays a crucial role in RLHF, yet it can be inefficient at the high-level due to limited samples (often only one per episode). In the operation of the high-levels, we find it necessary to increase the exploratory ability. As shown in Figure 8, reward learning becomes challenging without MaxEnt (maximum entropy). We compare RND and MaxEnt in terms of exploratory at high-levels. After conducting experiments displayed in Figure 10, we discovered that MaxEnt outperforms RND. While RND promotes the exploration of novel states, MaxEnt focuses on improving action entropy. The combination of the reward model and DDC can lead to previously disregarded decisions being recognized as good decisions. RND can learn and incorporate these decisions previously, treating them as non-novel at late times. However, MaxEnt does not encounter this issue and as a result, it achieves significantly better performance compared to RND.
w/ DDC
w/o DDC
@@
Figure 10: Exploratory mechanism effects analysis in FetchPickAndPlace domain.
Analyzing the overlay effects of DDC and human feedback. Having investigated DDC’s impact, we now turn to assess the combined influence of human guidance and DDC on subgoal formulation. Figure 9 depicts the training phase subgoal distribution, both with and without DDC. To the right of the figure, DDC’s inactivity results in a fragmented learning interaction between levels. While the high-level, aided by human guidance, swiftly navigates to the goal, it neglects the low-level’s execution capabilities, leading to transient and inefficient guidance, evident in the predominance of red subgoals in most regions except the upper-right corner. In contrast, on the figure’s left, with DDC active, the high-level delineates a subgoal sequence starting from the lower right, proceeding to the lower left, then the upper left, and culminating at the upper right corner. This strategy, different from the scenario without DDC, significantly bolsters the efficiency of human guidance, contributing to MENTOR overall effectiveness.
Figure 11: DDC effects on reward model in the Four rooms domain. The heatmaps serve as visual representations of the models’ learning progress, where the distance model learns to accurately estimate the state to initial state distances and the reward model learns to assign values that incentivize the agent’s progression towards the final goal. The overlay reward heatmap captures the integrated effect of both the reward model and the distance model, as articulated in the adjustments of the high-level policy update detailed in Equation 9.
We also analyze the heatmap of the reward and distance model. The analysis of the heatmaps in Figure 11 shows complex interactions between the distance and reward models in the MENTOR framework. The distance model assesses subgoal difficulty. This overlap, calculated with $r_{\text{hf}}(s,g^{\text{sub}},g)+\alpha H(g^{\text{sub}},g,k)$, identifies areas optimizing rewards in alignment with the low-level policy’s capabilities. However, without DDC, the human feedback-based reward model is limited, indicating only high-reward areas without guiding the agent on how to reach them. This can also corroborate the subgoal distribution in Figure 9.
IV-D Impact of Human Feedback (RQ3)
This part is dedicated to examining the effects of human feedback on Model MENTOR. Firstly, we evaluate the influence of the frequency and quantity of human feedback on algorithmic performance. Following this, we examine the differences between real human labels and synthetic labels, explain the rationale for using synthetic labels in our experiments, and demonstrate the practical applicability of our algorithm to real-world contexts.
TABLE II: Cross-analysis on episodes (100 steps per episode) to the success of variables batch queries and query frequency. For each combination of variables, we ran 5 trials on different seeds and recorded the mean and variance.
@@ IV-D Impact of Human Feedback (RQ3)
Figure 12: On the top side, it records the rate of disagreement between human annotations and synthetic labels. There lies a 10% to 20% disparity between human-labeled labels and synthetic labels. On the bottom side, we present the accuracy rate on the training data of the reward model. Within the Four room domain, the reward model’s accuracy is 75%, and approximately 80% for FetchPush.
Comparing human collected labels and synthetic labels. In our study, we employ two methods for providing human guidance: synthetic labels generated through scripted techniques and human-generated labels.
We monitored the disagreement rates between authentic human labels and synthetic labels, as well as the accuracy of the reward models, as shown in Figure 12. Given that the starting and ending positions in the Four rooms domain are predetermined, we provided 10 labels every 25 episodes. Conversely, in FetchPush where positions are not fixed, we provided 50 labels every 50 episodes. The data reveals discrepancies between human and synthetic labels, suggesting that human labeling is susceptible to errors due to subjective factors. Furthermore, the RLHF method produced a model whose accuracy converged to approximately 80%, indicating that even the trained model has its uncertainties and potential inaccuracies. This is also supported by the difference between the heatmap of the trained reward model and the heat map of the oracle reward model in Figure 11. However, our algorithm demonstrates sufficient robustness to handle the inaccuracies inherent in RLHF. This implies that integrating RLHF into the high-level policy is a highly compatible approach; on the one hand, RLHF fulfills the reward requirements of the high-level, and on the other hand, the low-level policy remains unaffected by the inaccuracies of RLHF. During the algorithm testing, the low-level policy completes tasks independently, hence the final performance of the algorithm is not affected.
We compare the performance of human collected labels and synthetic labels in FetchPush and Four rooms domains. As illustrated in Figure 13, models trained with human labels exhibited comparable learning rates, with performance differences potentially attributed to noise in human feedback. After comparing the experiments with human collected labels to the experiments with non-feedback, we can find a significant improvement in the algorithm’s convergence speed and final performance. Therefore, we conclude that synthetic labels have a similar effect on improving the algorithm as real human labels. This implies that in other experiments, synthetic labels could potentially serve as a replacement for human collected labels.
Figure 13: Experiments for evaluating MENTOR with script-generated and human-collected labels on the FetchPush and Four rooms domain. In these experiments, we provided 10 labels every 25 episodes. The results for the script-generated labels and non-feedback scenarios were based on five random seeds to ensure robustness, while the human-collected label experiment relied on a single random seed.
IV-E Ablation Studies (RQ4)